OpenStudy (anonymous):

I'm having trouble finding the centroid of a triangle algebraically. I have the midpoints and I also have the median equations, I just don't know how to get their point of intersection.

OpenStudy (klimenkov):

Write these equations here.

OpenStudy (anonymous):

The vertices are A(4,14) B(14,7) and C(2,4). The equations are: AD: y=-17/8x+45/2 CF: y=13/14x+15/7 EB: y=-2/11x+105/11 AD, CF, and EB are the Medians. I need to find their point of concurrence

OpenStudy (anonymous):

if all of these lines intersect each other, you just need to solve the linear system with any two of these equation

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

so I could just do -17/8x+45/2=13/14x+15/7 and that would give me the x, and then I just solve and find the y?

OpenStudy (anonymous):

that's it!

OpenStudy (anonymous):

Thank you soo much, that's a lot simpler than I thought!

OpenStudy (anonymous):

then you can plug this x value in all of the three equations and you will see that will return the same y value

OpenStudy (klimenkov):

Actually, if you have the vertices, you do not need these equations. The centroid will be \(\frac{A+B+C}{3}\).

OpenStudy (anonymous):

My teacher actually won't let us use that equation, but thank you!

OpenStudy (anonymous):

That is right @klimenkov , that is the most used way to find the centroid